OpenStudy (anonymous):

SOMEBODY PLEASE Solve for x. Type your answer as an integer, like this: 13 log5x + log4 = 3 HINT: Notice that the logs are in base 10.

OpenStudy (anonymous):

logx +logy=log(xy) so notice you multiply (xy) this is a standard formula log5x+log4=log(5x*4)=log20x now this is given as=3 log20x =3 the base 10^3=20x x=10*10*10/20 =50